OpenStudy (acacia):

What is the equation of a parabola with a vertex at (0, 0) and a directrix x = -4?

OpenStudy (anonymous):



OpenStudy (anonymous):

the distance between the vertex \((0,0)\) and the directrix is 4 because it is open to the right it looks like \((y-k)^2=4p(x-h)\) and in this case \(h=k=0\) and \(p=4\) so you get \[y^2=12x

OpenStudy (anonymous):

whoa that was wrong \(y^2=4\times 4x\) or \(y^2=16x\)

OpenStudy (acacia):

\[x=\frac{ 1 }{ 16 }y ^{2} ?\] is the answer?

OpenStudy (anonymous):

yes, that would be the same